OpenStudy (anonymous):

solve 4x^2-15x=3 how do i do that

OpenStudy (anonymous):

sum = -15, and product = 4*3

OpenStudy (anonymous):

write it as 4x^2 - 15x -3 = 0 check to see if this factorises first

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

how did you figure that out?

OpenStudy (anonymous):

yes - you can use quadratic formula - it does not factorise

OpenStudy (anonymous):

[-b+-sqrt(b^2-4*a*c)]/(2*a)

OpenStudy (anonymous):

used quadratic formula to solve for x: your given is a=4 b=-15 and c=-3

OpenStudy (anonymous):

a=4 b=-15 c=-3

OpenStudy (anonymous):

this is x =( -b +- sqrt(b^2 - 4ac)) / 2a for ax^2 + bx + c = 0

OpenStudy (anonymous):

have you used this formula before jenni?

OpenStudy (anonymous):

no it was a quiz and i have never used this before so i was wanting to know how to do it so next time i will get it right

OpenStudy (anonymous):

i got (2x+5)(2x+3) as the simplified form is that what i am supposed to do

OpenStudy (anonymous):

wait i meant factored form

OpenStudy (anonymous):

ok well as chris said a=4 b=-15 and c=-3 so plugging these values into the formula x = -(-15) + sqrt((-15)^2 - 4* 4*-3) / 2*4 = (15 + sqrt(273)) / 8 or ( 15 - sqrt(273)) / 8 this comes to 3.94 or -0.19 two roots

OpenStudy (anonymous):

as what jimmy gave you a quadratic formula x =( -b +- sqrt(b^2 - 4ac)) / 2a , just substitute all the given. a=4 b=-15 c=-3

OpenStudy (anonymous):

ok im really confused

OpenStudy (anonymous):

what? thats a different equation altogether

OpenStudy (anonymous):

2x+5)(2x+3) = 0 right?

OpenStudy (anonymous):

the simplest way to solve is by using the quadratic formula"-)

OpenStudy (anonymous):

(2x-5)(2x+3) was what i factored out 4x^2-15-3=0

OpenStudy (anonymous):

no those are the factors of 4x^2 -4x -15 (2x+5)(2x+3) = 4x^2 + 16x + 15

OpenStudy (anonymous):

the equation you originally posted cant be factored that why we used the formula

OpenStudy (anonymous):

anyway - i must go

OpenStudy (radar):

The quadratic formula is:\[-b \pm \sqrt{b ^{2}-4ac}\over 2a\] In the case of the original formula 4x^2-15x=3, it would be rewritten to standard form giving you \[4x ^{2}-15x-3=0\] where a=4,b=-15, and c=-3

OpenStudy (radar):

Plug those vales into the quadratic formula and you will get the roots of the equation.

OpenStudy (radar):

\[15\pm \sqrt{15^{2}+48}\over 8\] Can you take it from there?

OpenStudy (radar):

\[15\pm \sqrt{273}\over8\] \[15\pm16.522712 \over 8\]

OpenStudy (radar):

3.94, -0.19